# Quantum Rainbow Codes

### v2

@@ III.4 Seams, stabilisers and logical operators
Consider a pair of $D$-dimensional colour codes joined at a single splittable seam, as in Fig. 6. In unjoined colour codes stabilisers can be assigned to either $x$/$z$-maximal or $x$/$z$-rainbow subgraphs as these two types of subgraph are equivalent, but at the seam where two colour code lattices have been joined this is no longer the case. What possible assignments of stabilisers to subgraphs exist at the seam, and how do these assignments affect the logical operator structure of the resulting code?
@@
Each of the four classes of codes defined in Table 1 describes a different assignment of stabilisers to the subgraphs at the seam and in Table 2 we summarise how logical operators of the original codes are modified in each case (based on Lemma 7 and Lemma 8). The first three are fairly straightforward but it is interesting to discuss the mixed case in more detail, as the exact assignment of stabilisers for this case has not yet been explicitly defined.
@@
  • $2$-rainbow subgraphs for all other colourings
@@
The effect of the stabiliser assignment described in Lemma 9 is that, for parallel $d_{j}$
@@
For the case of parallel $d_{j}$, recall that in a colour code, in order to deform a $Z$ logical supported on $c_{i}$ edges across a 2D subregion of the code we must take the product of this logical with all $Z$ stabilisers supported on $\{c_{i},c_{j}\}$-rainbow subgraphs within this region. Recalling the proof of Lemma 7 we can then see that if $\{c_{0},c_{D}\}$-rainbow subgraphs of the joined lattice are not supports of $Z$ stabilisers, $c_{0}$ and $c_{D}$-coloured logicals parallel to the seam cannot be equivalent.
For the case of normal $d_{j}$ recall that $c_{i}$-coloured $Z$ string operators in a colour code anticommute with $X$ stabilisers only when these stabiliers are supported on subgraphs not containing edges of colour $c_{i}$. By Lemma 8 we then have that only $\overline{Z}_{(c_{i},d_{j})}^{m}$ with $c_{i}=c_{0}$ or $c_{D}$ anticommute with $X$ stabilisers at the seam. $\overline{Z}_{c_{i},d_{j}}^{m}$ for all other $c_{i}$ are therefore logical operators in the joined code, whereas for $c_{i}=c_{0}$ or $c_{D}$ only products $\overline{Z}_{(c_{i},d_{j})}^{1}\times\overline{Z}_{(c_{i},d_{j})}^{2}$ are logical operators. ∎
@@
$[\overline{Z}_{(c_{i},d_{j})}^{m},\overline{X}_{(c_{i},d_{j})}^{n}]=\begin{cases}-1\textrm{ for }m=n\\ 1\textrm{ for }m\neq n\end{cases}$ (18)
where $X$ operators are supported on $(D-1)$-dimensional hypermembranes containing all colours of edge except $c_{i}$ and normal to direction $d_{j}$. If after a join we have $\overline{Z}_{(c_{i},d_{j})}^{1}=\overline{Z}_{(c_{i},d_{j})}^{2}$ then $\overline{X}_{(c_{i},d_{j})}^{1}$ and $\overline{X}_{(c_{i},d_{j})}^{2}$ do not individually have consistent commutation relations with this operator and so only $\overline{X}_{(c_{i},d_{j})}^{1}\times\overline{X}_{(c_{i},d_{j})}^{2}$ is a logical $X$ operator. Examples of all possible logical operator transformations at seam are shown in Fig. 8.
All results in this section generalise straightforwardly to the case of more than two colour codes joined at a splittable seam. In this case Lemma 7 and Lemma 8 hold for all pairs of colour codes meeting at this seam, and so e.g. for parallel $Z$ logicals in generic codes we have $\overline{Z}_{(c_{i},d_{j})}^{1}=\overline{Z}_{(c_{i},d_{j})}^{2}=\overline{Z}_{(c_{i},d_{j})}^{3}=...$, and for normal $Z$ logicals in anti-generic codes $\overline{Z}_{(c_{i},d_{j})}^{1},\overline{Z}_{(c_{i},d_{j})}^{2},\overline{Z}_{(c_{i},d_{j})}^{3},...$ all anticommute with $X$ stabilisers so that only $\overline{Z}_{(c_{i},d_{j})}^{1}\times\overline{Z}_{(c_{i},d_{j})}^{2}\times\overline{Z}_{(c_{i},d_{j})}^{3}\times...$ is a logical operator.
We also want to consider cases such as Fig. 7 b), which can be interpreted as a joining of either a pair or a triple of colour codes. While we know how colour code logicals transform at the seams in this case there is an ambiguity about which colour code logicals we should consider (logicals of the pair of codes or of the triple). If we label the two options for cycle graphs used in the product as
@@
(where $\times$ is a composition of cycles by taking the symmetric difference of edges) then we can label the corresponding sets of colour code lattices as $\{C_{1},C_{2},C_{3}\}$ and $\{C_{1\times 2},C_{3}\}$. If we then consider a generic rainbow code defined on the joined lattices we can see that, in the former case, we have $\overline{Z}_{(c_{i},d_{j})}^{1}$, $\overline{Z}_{(c_{i},d_{j})}^{2}$ and $\overline{Z}_{(c_{i},d_{j})}^{3}$ all as independent logical operators for $d_{j}$ normal to both seams and for any $c_{i}$. On the other hand, in the latter case we have just $\overline{Z}_{(c_{i},d_{j})}^{1\times 2}$ and $\overline{Z}_{(c_{i},d_{j})}^{3}$. This is therefore not a complete basis, and in fact $\overline{Z}_{(c_{i},d_{j})}^{1\times 2}=\overline{Z}_{(c_{i},d_{j})}^{1}\times\overline{Z}_{(c_{i},d_{j})}^{2}$ as can be checked using Fig. 7 b). We therefore conclude that in order to identify a complete basis for colour code logical operators in such a code we must identify a complete cycle basis for the graphs used as input to the product and consider the colour code lattices obtained from these cycles. In contrast, the number of physical qubits in this code is equal to the number of physical qubits in $C_{1\times 2}\cup C_{3}$ (compare Fig. 7 b) to Fig. 6 a)), which is fewer than the number of qubits we would have in $C_{1}\cup C_{2}\cup C_{3}$. These codes thus have the potential for improved encoding rates relative to disjoint collections of colour codes.
@@ III.5 Hypergraph product rainbow codes
$\{g_{i}\Leftarrow g_{j}\leavevmode\nobreak\ \forall\leavevmode\nobreak\ (g_{i},g_{j})\in P_{k}\}:\mathcal{G}_{k}^{\prime}$ (20)
@@
Each $\mathcal{G}_{k}^{\prime}$ is a collection of disjoint cycles and so each connected component of $_{k}^{D}\mathcal{G}_{k}^{\prime}$ is a $D$-dimensional hypercubic lattice on a $D$-dimensional torus and the flag graph of each of these components is a $D$-dimensional colour code lattice by Lemma 5. If we label the elements of the cycle basis of each $\mathcal{G}_{k}^{\prime}$ as $\mathcal{O}_{a}^{k}$ (for $1\leq a\leq n_{c}^{k}$) then each colour code lattice is indexed by $(\mathcal{O}_{a_{1}}^{1},\mathcal{O}_{a_{2}}^{2},...,\mathcal{O}_{a_{D}}^{D})$ and these can be interpreted as coordinates, giving a natural arrangement of these lattices in a $D$-dimensional grid e.g. Fig. 9. Each of these lattices defines a ($x=D$ and $z=2$) colour code with logical $Z$ operators which we can index as $(\mathcal{O}_{a_{1}}^{1},\mathcal{O}_{a_{2}}^{2},...,\mathcal{O}_{a_{D}}^{D};c_{i},d_{j})$ where $c_{i}$ and $d_{j}$ are colour and direction as before, with $d_{k}$ being the direction in the grid associated with the cycles of $\mathcal{G}_{k}^{\prime}$. For a given code and fixed $d_{j}$ only $D$ logicals of each colour are independent, so we have $D^{2}$ logical qubits per code.
@@
The gluings $\{g_{i}\Leftarrow g_{j}\ \leavevmode\nobreak\ \forall\leavevmode\nobreak\ (g_{i},g_{j})\in P_{k}\}$ join all the disjoint cycles of $\mathcal{G}_{k}^{\prime}$ into a single connected component and so join together all colour code lattices lying along each $k$-directional line of the grid (e.g. $\{g_{i}\Leftarrow g_{j}\ \leavevmode\nobreak\ \forall\leavevmode\nobreak\ (g_{i},g_{j})\in P_{1}\}$ join the lattices of each row of Fig. 9 while $\{g_{i}\Leftarrow g_{j}\ \leavevmode\nobreak\ \forall\leavevmode\nobreak\ (g_{i},g_{j})\in P_{2}\}$ join the lattices of each column). Because the $\mathcal{G}_{k}$ contain only even-degree vertices the resulting seams are all splittable seams by Lemma 6, and we know how colour code logical operators transform at each of these seams by Lemma 7 and Lemma 8. This allows us to understand the resulting structure of these operators at the level of the grid. Explicitly, for each of our previously defined classes of codes, we have the following:
@@
Generic: In these codes, for any choice of cycle $\mathcal{O}_{a_{k}}^{k}$, all logical $Z$ operators of the form $(\mathcal{O}_{a_{1}}^{1},...,\mathcal{O}_{a_{k}}^{k},...,\mathcal{O}_{a_{D}}^{D};c_{i},d_{k})$ for fixed $c_{i}$ are equivalent, essentially giving one stringlike logical of each colour associated to each $\mathcal{O}_{a_{k}}^{k}$ (red lines in Fig. 9). Recalling that in each colour code only $D$ colours of logical are independent for fixed $d_{j}$ this gives a number of independent logical operators
@@
Anti-generic: In these codes, for any choice of cycles $\mathcal{O}_{a_{1}}^{1},...,\mathcal{O}_{a_{k-1}}^{k-1},\mathcal{O}_{a_{k+1}}^{k_{1}},...,\mathcal{O}_{a_{D}}^{D}$, all logical $Z$ operators of the form $(\mathcal{O}_{a_{1}}^{1},...,\mathcal{O}_{a_{k}}^{k},...,\mathcal{O}_{a_{D}}^{D};c_{i},d_{k})$ are merged into a single logical operator, giving one logical of each colour for each $k$-directional line of the grid (blue lines in Fig. 9). The number of independent logicals in this case is then
$n_{L}=\sum_{i}\prod_{j\neq i}Dn_{c}^{i}$ (25)
@@
and distance linear in the girth of the input graphs. However, we also note that for fixed input graphs the distance of a mixed code defined on the resulting flag graph will be twice the distance of the generic code defined on this same graph (as long as all $\mathcal{G}_{k}^{\prime}$ contain more than one connected component). This is due to the fact that the $c_{0}$- and $c_{D}$-coloured logical $Z$ operators of a colour code have half the weight of the logicals of all other colours, but in the mixed code the $c_{0}$/$c_{D}$ logicals must be extended across multiple component colour codes and so the new lowest-weight logicals will be those of other colours.
@@ IV Logical gates of HGP rainbow codes
The following are necessary and sufficient conditions for a quantum CSS code to possess a transversal non-Clifford gate implemented by application of physical $T$/$T^{\dagger}$ to a bipartition of the qubits.
@@
Due to Proposition B.3 of CSSLO , $W$ is a diagonal logical operator if and only if the group commutator $[[X(\mathbf{x_{i}}),W]]$ is a logical identity for all rows $\mathbf{x_{i}}$ of ${S_{X}}$. Calculating the group commutator using the identity in Table 4 of XP :
@@
The first condition is equivalent to 4 in the Lemma. Turning to the second condition, $\mathbf{x}-2\mathbf{ax}_{i}\in\braket{M_{Z}}$ if and only if the group commutator $[[X(\mathbf{x}_{j}),S(\mathbf{x}_{i}-2\mathbf{ax}_{i})]]$ is a logical identity for all $\mathbf{x}_{j}\in\braket{S_{X},L_{X}}$. Due to Proposition E.14 of XP , it is sufficient to consider only $\mathbf{x}_{j}$ which are rows and sums of pairs of rows from the matrix $\begin{pmatrix}S_{X}\\ L_{X}\end{pmatrix}$. Calculating the group commutator:
@@
Pin: By Lemma 2 the intersections of $X$ stabilisers are supports of $Z$ stabilisers, satisfying 1. However, because $x$-rainbow subgraphs support $X$ logical operators, by Lemma 3 we have that the intersections of $X$ logicals and $X$ stabilisers can be $Z$ logicals and so 2 is not generally satsified in these codes.
Generic: By Lemma 2 the intersections of $X$ stabilisers are $2$-maximal subgraphs, but the supports of $Z$ stabilisers are $2$-rainbow subgraphs. Thus 1 is only satisfied if we can always decompose these $2$-maximal subgraphs into $2$-rainbow subgraphs by deleting some of the edges. Fortunately this is true in our case because all seams are splittable seams and by deleting all edges created by the gluing we recover a disjoint collection of colour code lattices. This same property also allows us to split $3$-maximal subgraphs into disjoint collections of $3$-rainbow subgraphs which are $X$ stabilisers of the component colour codes. The logical operators of the rainbow code are also just logical operators of the component colour codes and, since we know 2 and 3 are satisfied in these colour codes, they must also be satisfied in the rainbow code.

### v3

@@ III.4 Seams, stabilisers and logical operators
Consider a pair of $D$-dimensional colour codes joined at a single splittable seam, as in Fig.˜6. In unjoined colour codes stabilisers can be assigned to either xx/$z$-maximal or xx/$z$-rainbow subgraphs as these two types of subgraph are equivalent, but at the seam where two colour code lattices have been joined this is no longer the case. What possible assignments of stabilisers to subgraphs exist at the seam, and how do these assignments affect the logical operator structure of the resulting code?
@@
Each of the four classes of codes defined in Table˜1 describes a different assignment of stabilisers to the subgraphs at the seam and in Table˜2 we summarise how logical operators of the original codes are modified in each case (based on Lemma˜8 and Lemma˜9). The first three are fairly straightforward but it is interesting to discuss the mixed case in more detail, as the exact assignment of stabilisers for this case has not yet been explicitly defined.
@@
  • 22-rainbow subgraphs for all other colourings
@@
The effect of the stabiliser assignment described in Lemma˜10 is that, for parallel $d_{j}$
@@
For the case of parallel $d_{j}$, recall that in a colour code, in order to deform a $Z$ logical supported on $c_{i}$ edges across a 2D subregion of the code we must take the product of this logical with all $Z$ stabilisers supported on $\{c_{i},c_{j}\}$-rainbow subgraphs within this region. Recalling the proof of Lemma˜8 we can then see that if $\{c_{0},c_{D}\}$-rainbow subgraphs of the joined lattice are not supports of $Z$ stabilisers, $c_{0}$ and $c_{D}$-coloured logicals parallel to the seam cannot be equivalent.
For the case of normal $d_{j}$ recall that $c_{i}$-coloured $Z$ string operators in a colour code anticommute with $X$ stabilisers only when these stabiliers are supported on subgraphs not containing edges of colour $c_{i}$. By Lemma˜9 we then have that only $\overline{Z}_{(c_{i},d_{j})}^{m}$ with $c_{i}=c_{0}$ or $c_{D}$ anticommute with $X$ stabilisers at the seam. $\overline{Z}_{c_{i},d_{j}}^{m}$ for all other $c_{i}$ are therefore logical operators in the joined code, whereas for $c_{i}=c_{0}$ or $c_{D}$ only products $\overline{Z}_{(c_{i},d_{j})}^{1}\times\overline{Z}_{(c_{i},d_{j})}^{2}$ are logical operators. ∎
@@
where $X$ operators are supported on (D1)(D-1)-dimensional hypermembranes containing all colours of edge except $c_{i}$ and normal to direction $d_{j}$. If after a join we have $\overline{Z}_{(c_{i},d_{j})}^{1}=\overline{Z}_{(c_{i},d_{j})}^{2}$ then $\overline{X}_{(c_{i},d_{j})}^{1}$ and $\overline{X}_{(c_{i},d_{j})}^{2}$ do not individually have consistent commutation relations with this operator and so only $\overline{X}_{(c_{i},d_{j})}^{1}\times\overline{X}_{(c_{i},d_{j})}^{2}$ is a logical $X$ operator. Examples of all possible logical operator transformations at seam are shown in Fig.˜8.
All results in this section generalise straightforwardly to the case of more than two colour codes joined at a splittable seam. In this case Lemma˜8 and Lemma˜9 hold for all pairs of colour codes meeting at this seam, and so e.g. for parallel $Z$ logicals in generic codes we have $\overline{Z}_{(c_{i},d_{j})}^{1}=\overline{Z}_{(c_{i},d_{j})}^{2}=\overline{Z}_{(c_{i},d_{j})}^{3}=$..., and for normal $Z$ logicals in anti-generic codes $\overline{Z}_{(c_{i},d_{j})}^{1},\overline{Z}_{(c_{i},d_{j})}^{2},\overline{Z}_{(c_{i},d_{j})}^{3},$... all anticommute with $X$ stabilisers so that only $\overline{Z}_{(c_{i},d_{j})}^{1}\times\overline{Z}_{(c_{i},d_{j})}^{2}\times\overline{Z}_{(c_{i},d_{j})}^{3}\times$... is a logical operator.
We also want to consider cases such as Fig.˜7 b), which can be interpreted as a joining of either a pair or a triple of colour codes. While we know how colour code logicals transform at the seams in this case there is an ambiguity about which colour code logicals we should consider (logicals of the pair of codes or of the triple). If we label the two options for cycle graphs used in the product as
[…]
(where $\times$ is a composition of cycles by taking the symmetric difference of edges) then we can label the corresponding sets of colour code lattices as $\{C_{1},C_{2},C_{3}\}$ and $\{C_{1\times 2},C_{3}\}$. If we then consider a generic rainbow code defined on the joined lattices we can see that, in the former case, we have $\overline{Z}_{(c_{i},d_{j})}^{1}$, $\overline{Z}_{(c_{i},d_{j})}^{2}$ and $\overline{Z}_{(c_{i},d_{j})}^{3}$ all as independent logical operators for $d_{j}$ normal to both seams and for any $c_{i}$. On the other hand, in the latter case we have just $\overline{Z}_{(c_{i},d_{j})}^{1\times 2}$ and $\overline{Z}_{(c_{i},d_{j})}^{3}$. This is therefore not a complete basis, and in fact $\overline{Z}_{(c_{i},d_{j})}^{1\times 2}=\overline{Z}_{(c_{i},d_{j})}^{1}\times\overline{Z}_{(c_{i},d_{j})}^{2}$ as can be checked using Fig.˜7 b). We therefore conclude that in order to identify a complete basis for colour code logical operators in such a code we must identify a complete cycle basis for the graphs used as input to the product and consider the colour code lattices obtained from these cycles. In contrast, the number of physical qubits in this code is equal to the number of physical qubits in $C_{1\times 2}\cup C_{3}$ (compare Fig.˜7 b) to Fig.˜6 a)), which is fewer than the number of qubits we would have in $C_{1}\cup C_{2}\cup C_{3}$. These codes thus have the potential for improved encoding rates relative to disjoint collections of colour codes.
@@ III.5 Hypergraph product rainbow codes
$\{g_{i}\Leftarrow g_{j}\penalty 10000\ \forall\penalty 10000\ (g_{i},g_{j})\in P_{k}\}:\mathcal{G}_{k}^{\prime}$ (20)
@@
[…]
@@
The gluings $\{g_{i}\Leftarrow g_{j}\ \penalty 10000\ \forall\penalty 10000\ (g_{i},g_{j})\in P_{k}\}$ join all the disjoint cycles of $\mathcal{G}_{k}^{\prime}$ into a single connected component and so join together all colour code lattices lying along each $k$-directional line of the grid (e.g. $\{g_{i}\Leftarrow g_{j}\ \penalty 10000\ \forall\penalty 10000\ (g_{i},g_{j})\in P_{1}\}$ join the lattices of each row of Fig.˜9 while $\{g_{i}\Leftarrow g_{j}\ \penalty 10000\ \forall\penalty 10000\ (g_{i},g_{j})\in P_{2}\}$ join the lattices of each column). Because the $\mathcal{G}_{k}$ contain only even-degree vertices the resulting seams are all splittable seams by Lemma˜7, and we know how colour code logical operators transform at each of these seams by Lemma˜8 and Lemma˜9. This allows us to understand the resulting structure of these operators at the level of the grid. Explicitly, for each of our previously defined classes of codes, we have the following:
@@
Generic: In these codes, for any choice of cycle $\mathcal{O}_{a_{k}}^{k}$, all logical $Z$ operators of the form $(\mathcal{O}_{a_{1}}^{1},...,\mathcal{O}_{a_{k}}^{k},...,\mathcal{O}_{a_{D}}^{D};c_{i},d_{k})$ for fixed $c_{i}$ are equivalent, essentially giving one stringlike logical of each colour associated to each $\mathcal{O}_{a_{k}}^{k}$ (red lines in Fig.˜9). Recalling that in each colour code only $D$ colours of logical are independent for fixed $d_{j}$ this gives a number of independent logical operators
@@
Anti-generic: In these codes, for any choice of cycles $\mathcal{O}_{a_{1}}^{1},...,\mathcal{O}_{a_{k-1}}^{k-1},\mathcal{O}_{a_{k+1}}^{k_{1}},...,\mathcal{O}_{a_{D}}^{D}$, all logical $Z$ operators of the form $(\mathcal{O}_{a_{1}}^{1},...,\mathcal{O}_{a_{k}}^{k},...,\mathcal{O}_{a_{D}}^{D};c_{i},d_{k})$ are merged into a single logical operator, giving one logical of each colour for each $k$-directional line of the grid (blue lines in Fig.˜9). The number of independent logicals in this case is then
$n_{L}=\sum_{i}\prod_{j\neq i}Dn_{c}^{j}$ (25)
@@
and distance linear in the girth of the input graphs. However, we also note that for fixed input graphs the distance of a mixed code defined on the resulting flag graph will be twice the distance of the generic code defined on this same graph (as long as all $\mathcal{G}_{k}^{\prime}$ contain more than one connected component). This is due to the fact that the $c_{0}$- and $c_{D}$-coloured logical $Z$ operators of a colour code have half the weight of the logicals of all other colours, but in the mixed code the $c_{0}$$c_{D}$ logicals must be extended across multiple component colour codes and so the new lowest-weight logicals will be those of other colours.
All of the families of codes discussed above are LDPC. This follows from the fact that the weights of the stabiliser generators are upper-bounded by the size of the largest $D$-maximal subgraph, which, in the language of pin codes, is the size of the largest $1$-pinned set. It was shown in vuillot2022quantum that pin codes on simplicial complexes obtained from the hypergraph product of classical LDPC codes are themselves LDPC, and so this is also true of rainbow codes defined on the same complexes.
@@ IV Logical gates of HGP rainbow codes
The following are necessary and sufficient conditions for a quantum CSS code to possess a transversal non-Clifford gate implemented by application of physical $TT^{\dagger}$ to a bipartition of the qubits.
@@
Due to Proposition B.3 of CSSLO, $W$ is a diagonal logical operator if and only if the group commutator $X(\mathbf{x_{i}}),W]]$ is a logical identity for all rows $\mathbf{x_{i}}$ of $S_{X}$}. Calculating the group commutator using the identity in Table 4 of XP:
@@
The first condition is equivalent to 4 in the Lemma. Turning to the second condition, $\mathbf{x}-2\mathbf{ax}_{i}\in\braket{M_{Z}}$ if and only if the group commutator $X(\mathbf{x}_{j}),S(\mathbf{x}_{i}-2\mathbf{ax}_{i})]]$ is a logical identity for all $\mathbf{x}_{j}\in\braket{S_{X},L_{X}}$. Due to Proposition E.14 of XP, it is sufficient to consider only $\mathbf{x}_{j}$ which are rows and sums of pairs of rows from the matrix $\begin{pmatrix}S_{X}\\ L_{X}\end{pmatrix}$. Calculating the group commutator:
@@
Pin: By Lemma˜3 the intersections of $X$ stabilisers are supports of $Z$ stabilisers, satisfying 1. However, because $x$-rainbow subgraphs support $X$ logical operators, by Lemma˜4 we have that the intersections of $X$ logicals and $X$ stabilisers can be $Z$ logicals and so 2 is not generally satsified in these codes.
Generic: By Lemma˜3 the intersections of $X$ stabilisers are $2$-maximal subgraphs, but the supports of $Z$ stabilisers are $2$-rainbow subgraphs. Thus 1 is only satisfied if we can always decompose these $2$-maximal subgraphs into $2$-rainbow subgraphs by deleting some of the edges. Fortunately this is true in our case because all seams are splittable seams and by deleting all edges created by the gluing we recover a disjoint collection of colour code lattices. This same property also allows us to split $3$-maximal subgraphs into disjoint collections of $3$-rainbow subgraphs which are $X$ stabilisers of the component colour codes. The logical operators of the rainbow code are also just logical operators of the component colour codes and, since we know 2 and 3 are satisfied in these colour codes, they must also be satisfied in the rainbow code.